### v1

# MorphoSeg: An Uncertainty-Aware Deep Learning Method for Biomedical Segmentation of Complex Cellular Morphologies

Abstract
Deep learning has revolutionized medical and biological imaging, particularly in segmentation tasks. However, segmenting biological cells remains challenging due to the high variability and complexity of cell shapes. Addressing this challenge requires high-quality datasets that accurately represent the diverse morphologies found in biological cells. Existing cell segmentation datasets are often limited by their focus on regular and uniform shapes. In this paper, we introduce a novel benchmark dataset of Ntera-2 (NT2) cells, a pluripotent carcinoma cell line, exhibiting diverse morphologies across multiple stages of differentiation, capturing the intricate and heterogeneous cellular structures that complicate segmentation tasks. To address these challenges, we propose an uncertainty-aware deep learning framework for complex cellular morphology segmentation (MorphoSeg) by incorporating sampling of virtual outliers from low-likelihood regions during training. Our comprehensive experimental evaluations against state-of-the-art baselines demonstrate that MorphoSeg significantly enhances segmentation accuracy, achieving up to a 7.74% increase in the Dice Similarity Coefficient (DSC) and a 28.36% reduction in the Hausdorff Distance. These findings highlight the effectiveness of our dataset and methodology in advancing cell segmentation capabilities, especially for complex and variable cell morphologies. The dataset and source code is publicly available at https://github.com/RanchoGoose/MorphoSeg.
keywords:
Biomedical Segmentation, Cell Segmentation, Machine Learning, Deep Learning, Ntera-2 Cells, Data Repository, Complex Cell Shapes, Vision Transformer
journal: Neurocomputing Journal
[…]
[1]organization=School of Electrical and Electronic Enigeering, The University of Sheffield, city=Sheffield, country=United Kingdom
[…]
[2]organization=School of Biosciences, The University of Sheffield, city=Sheffield, country=United Kingdom
1 Introduction
The advent of deep learning has significantly advanced the field of medical imaging, particularly in segmentation tasks. While substantial progress has been made in segmenting well-defined objects such as organs in computed tomography (CT) scans [1, 2, 3, 4], cell segmentation presents unique challenges due to the complexity and variability of cell shapes [5, 6, 7, 8]. Existing datasets for cell segmentation [9, 10, 11, 12] often fall short in covering the full spectrum of cell types, hindering the effective training of deep learning models for diverse research applications.
In the broader context of medical image analysis, cell segmentation is crucial yet challenging, particularly due to the high variability in cell stages and the resolutions of medical images. Current state-of-the-art (SOTA) deep learning models, including advanced architectures such as Convolutional Neural Networks (CNNs) [6, 13] and Transformers [1, 3], typically require large, well-annotated datasets, which are resource-intensive to acquire. Beyond focusing on model algorithms, some researchers have shifted towards developing methods for rapid, easy data annotation or advancing weak, semi, or unsupervised learning algorithms to address data insufficiency [14, 15]. However, irregular cell shapes and the need for multi-stage observation remain significant challenges.
In this paper, we present a novel benchmark dataset featuring Ntera-2 cells (NT2) cells [16]. NT2 is a pluripotent carcinoma cell line known for its ability to differentiate into neuro-ectodermal lineages [17]. NT2 cells have previously been shown to transform into neurons, after exposure to all-trans-retinoic acid [18] (ATRA). During differentiation, NT2 cells exhibit a range of heterogeneous morphologies, as shown in Figure 1, which poses challenges for reliable segmentation. The dataset includes images of NT2 cells at numerous stages of differentiation, spanning from day 7 to day 12, and captures the diverse and complex cell morphologies formed during this period. The complexities can make it difficult to accurately annotate cells, with some cellular features barely discernible even to trained experts. Furthermore, the variability in shape across different stages complicates the segmentation process, necessitating robust deep-learning solutions capable of adapting to such heterogeneity.
Figure 1: Undifferentiated and Differentiating NT2 Cell Example Images From the Training Set. (A) Undifferentiated Ntera-2 cells. (B) shows the original bright-field images of cells exposed to $10^{-5}$ M ATRA every 48 hours for 2 weeks. (C) displays the mask created from manual annotations. The images highlight the diverse data set used for training.
To address the aforementioned challenges, we propose a novel unknown-aware enhancement algorithm based on the SOTA medical segmentation framework TransUNet [1]. Our approach involves sampling virtual outliers from the low-likelihood regions of the class-conditional distributions during the training stage to improve the decision boundaries of the segmented objects. This strategy enhances the robustness of the deep learning models, making them effective solutions in such complex scenarios. By incorporating uncertainties into the training objective, our proposed deep learning model achieves strong segmentation of cells with random shapes. When compared against several benchmarks, we demonstrate significant improvements with our proposed dataset: the mean Dice Similarity Coefficient (DSC) increased from 80.35% to 86.57% and the mean Hausdorff Distance (HD95) decreased from 21.98% to 15.75%. These results confirm that our approach not only mitigates data scarcity but also enhances the model’s generalization capabilities across varied cell morphologies. We have provided the full code and pre-trained models with simple inference instructions to facilitate easy implementation and further research by field researchers.
The key contributions of this paper are:
  1. 1.
    Development of a pluripotent carcinoma cell repository, specifically curated to support the development and validation of advanced biological cell segmentation models.
  2. 2.
    Proposal of a practical and scalable framework for cell segmentation that leverages synthetic outliers and data augmentation to significantly enhance the performance of the SOTA cell segmentation deep-learning models, particularly in scenarios typical of limited data availability.
  3. 3.
    Comprehensive quantitative and qualitative analyses highlighting that the proposed model performs favourably compared to existing techniques.
The remainder of this paper is organized as follows: Section 2 reviews the literature on deep learning methods for biological cell segmentation, and outlier generation methods. Section 3 details our proposed model’s methodology and framework. Section 4 elaborates on the dataset and preprocessing techniques. Section 5 presents experimental setup, results, analyses, and ablation studies, as well as inference and visualization techniques. Finally, Section 6 concludes the paper, summarizing key findings and future research directions.
2 Related Work
2.1 Medical and Biological Image Segmentation.
In recent years, deep learning techniques in computer vision have achieved promising performance in medical image segmentation. Encoder-decoder-based CNN structures have been widely applied, with U-Net and its variants significantly advancing the field [2, 13, 19, 20, 21]. Building upon this structure, UNet++ incorporates densely connected links to enhance the fusion of multi-scale features, improving the analysis of medical images across CT, MRI, and EM modalities [22]. Furthermore, nnU-Net, a self-configuring U-Net, has shown remarkable performance across various segmentation tasks with minimal manual parameter tuning [23, 24]. In terms of segmenting cell structures, the Cellpose model utilizes a two-headed output architecture to predict cell probabilities and spatial organization, effectively handling cells of regular shapes [6]. However, its performance is limited for irregularly shaped cells, as often encountered in advanced disease stages.
Semantic segmentation is pivotal for clinical tasks such as disease detection, differential diagnosis, survival prediction, therapy planning, and treatment response assessment [25, 26, 4, 27, 28, 29]. Besides deep learning methods based on CNN structures, attention mechanisms [30] are extensively employed in semantic segmentation to prioritize salient features. The integration of Vision Transformers (ViTs) [31] into medical segmentation networks has enhanced performance by leveraging self-attention mechanisms [32, 33, 34]. Some research focuses on combining the strengths of transformers with traditional CNN architectures, creating hybrid encoder structures. TransUNet combines CNNs for spatial information extraction with transformers for capturing long-range dependencies, featuring a U-Net structure with self-attention mechanisms to ensure comprehensive image analysis [1]. This integration makes TransUNet particularly effective for detailed medical and biological image segmentation, where preserving spatial hierarchies is crucial. Other research includes employing channel attention to capture boundary-aware features for enhancing polyp segmentation [35]. CellViT replaces the CNN encoder with a transformer block in the U-Net architecture [36]. The recent SAMed model extends the capabilities of the Segment Anything Model (SAM) [37] to medical imaging by employing a Low-rank Adaptation (LoRA) fine-tuning strategy to adapt the large-scale image encoder for medical contexts effectively [3]. This method illustrates the advantage of using extensive pre-trained models for specialized medical segmentation tasks. However, SAM is designed for multi-class segmentation, posing limitations for binary segmentation scenarios.
2.2 Virtual Outliers Synthesis
Outliers in data mining and statistics are typically considered representations of abnormalities, discordants, deviants, or anomalies [38, 39]. In the context of deep learning, outliers are often viewed as anomalies compared to inliers, with multiple deep learning methods developed to detect these outliers [40, 39]. Unlike traditional approaches that focus on outlier detection, recent works in computer vision have explored synthesizing outliers to enhance the accuracy and robustness of deep learning models, particularly in scenarios where datasets are insufficient, such as the medical image classification [39].
Further research has leveraged synthetic data to quantify uncertainty, primarily for out-of-distribution (OOD) detection [41, 42, 4]. OOD detection aims to identify conditions unrepresented in the training dataset [43]. In the realm of OOD detection, the challenge is to detect OOD samples; in our case, these include unseen cells in annotations that are difficult for precise human annotation and cells at transitional stages that can be considered OOD. The subtle differences between foreground elements in medical images turn OOD detection into a near-OOD problem, where traditional methods often fall short [44, 45, 46, 47]. In our research, we treat segmentation targets, specifically NT2 cells at various stages, as OOD data. Virtual outlier synthesis has proven effective for OOD detection [41], inspiring us to adapt this approach to enhance the robustness and accuracy of our model for cell segmentation.
Synthesis methods for medical and biological image segmentation remain relatively underexplored. Few synthesis methods have been applied to enhance the training of deep learning models in this area [48, 49, 50]. Our proposed method dynamically estimates class-conditional Gaussian distributions during training, shaping the uncertainty surface over time with our specially formulated loss.
Figure 2: The overall framework of the proposed method, following the architecture of TransUNet [1]. The process for virtual outlier sampling is depicted with red dotted lines. In this framework, we model the feature representations as class-conditional Gaussians and sample virtual outliers $v$ from regions of low likelihood. These virtual outliers contribute to generating the uncertainty loss for regularization. The synthetic segmentation map, incorporating the uncertainty estimation branch $\mathcal{L}_{\text{uncertainty}}$, is trained concurrently with the mixed segmentation loss $\mathcal{L}_{\text{seg}}$ in Eq. (12).
3 Methodology
The overall framework of our approach is depicted in Figure 2. We utilize the TransUNet [1] architecture as the foundational framework for our method. Initially, high-resolution images are subdivided into small patches to align with the input constraints of the vision transformer and enhance the model’s ability to capture cellular details at the pixel level (details in Section 4.2). For the original TransUNet, there are two types of encoders for feature extraction: a hybrid encoder combining a pre-trained CNN and a pure Transformer encoder. We present the hybrid structure in Figure 2 as it provides better results (shown in Section 5.3). It is noteworthy that our method is scalable to different deep learning frameworks since the sampling method works on the raw segmentation predictions (details in Section 3.2). The rest of this section is orgnized as follows: Section 3.1 provides preliminary information and an overview of the TransUNet framework for image segmentation tasks. In Section 3.2, we present our robust cell segmentation enhancement via virtual outlier sampling. Section 3.3 introduces the overall uncertainty-aware training objective for MorphoSeg.
3.1 TransUNet for Cell Segmentation
Given an image $\bm{\mathrm{x}}\in\mathbb{R}^{H\times W\times C}$ with a spatial resolution of $H\times W$ and $C$ channels, our objective is to predict the corresponding pixel-wise label map of the same dimensions. Traditionally, this is achieved by training a CNN (e.g., U-Net) that encodes the image into high-level feature representations and subsequently decodes these representations back to the original spatial resolution. To enhance segmentation accuracy, particularly for cell segmentation, we incorporate a Vision Transformer with self-attention mechanisms for instance segmentation without losing fine-grained information.
Image Sequentialization. For a pure Transformer encoder structure, following [31], we first perform tokenization by reshaping the input $\bm{\mathrm{x}}$ into a sequence of flattened 2D patches $\bm{\mathrm{x}}^{i}_{p}\in\mathbb{R}^{P^{2}\cdot C}|i=1,..,N\}$, where each patch is of size $P\times P$ and $N=\frac{HW}{P^{2}}$ is the number of image patches (i.e., the input sequence length).
Patch Embedding. We map the vectorized patches $\bm{\mathrm{x}}_{p}$ into a latent $D$-dimensional embedding space using a trainable linear projection. To encode the patch spatial information, we learn specific position embeddings which are added to the patch embeddings to retain positional information as follows:
$\bm{\mathrm{z}}_{0}$ $=[\bm{\mathrm{x}}^{1}_{p}\bm{\mathrm{E}};\,\bm{\mathrm{x}}^{2}_{p}\bm{\mathrm{E}};\cdots;\,\bm{\mathrm{x}}^{N}_{p}\bm{\mathrm{E}}]+\bm{\mathrm{E}}_{pos},$ (1)
where $\bm{\mathrm{E}}\in\mathbb{R}^{(P^{2}\cdot C)\times D}$ is the patch embedding projection, and $\bm{\mathrm{E}}_{pos}\in\mathbb{R}^{N\times D}$ denotes the position embedding.
The concatenated sequence $\mathbf{T}$ is processed through the Transformer encoder layers to yield the final representations. Following the the architecture of the vanilla ViT [31], we employ multi-headed self-attention (MSA) [30] (Eq. 3) and multi-layer perceptron (MLP) blocks (Eq. 4). Layer normalization (LN) is applied before each block (Eq. 5), as described in the following equations:
$\mathbf{Z}_{0}$ $=\mathbf{T},$ (2)
$\mathbf{Z}^{\prime}_{\ell}$ $=\text{MSA}(\text{LN}(\mathbf{Z}_{\ell-1}))+\mathbf{Z}_{\ell-1},\quad\ell=1\ldots L,$ (3)
$\mathbf{Z}_{\ell}$ $=\text{MLP}(\text{LN}(\mathbf{Z}^{\prime}_{\ell}))+\mathbf{Z}^{\prime}_{\ell},\quad\ell=1\ldots L,$ (4)
$\mathbf{Y}$ $=\text{LN}(\mathbf{Z}_{L}[0]),$ (5)
where $L$ denotes the number of Transformer layers, and $\mathbf{Z}_{L}[0]$ represents the final layer’s class token representation. The output $\mathbf{Y}\in\mathbb{R}^{D}$ serves as the input to a classifier head for the task at hand. For segmentation, an effective approach is to upsample the encoded feature representation $\bm{\mathrm{z}}_{L}\in\mathbb{R}^{\frac{HW}{P^{2}}\times D}$ to the full resolution. This involves reshaping the encoded feature from $\frac{HW}{P^{2}}$ to $\frac{H}{P}\times\frac{W}{P}$, applying a $1\times 1$ convolution to reduce the channel size to the number of classes, and then bilinearly upsampling the feature map to the full resolution $H\times W$ for the final segmentation outcome.
Due to the limited dimensions of ViT input tokens, input image patch sizes $\frac{H}{P}\times\frac{W}{P}$ are often much smaller than the original image resolution $H\times W$ in high-resolution medical images, leading to a loss of low-level details (e.g., cell shape and boundary). TransUNet addresses this by using a CNN-Transformer hybrid model, where a CNN first extracts a feature map from the input. Patch embedding is then applied to $1\times 1$ patches from the CNN feature map instead of raw images, ensuring better retention of details. This method’s efficacy is demonstrated in the comparative analysis of different encoders presented in Table 1 in Section 5.3.
Cascaded Upsampler. Instead of naive upsampling in a ViT, a Cascaded Upsampler (CUP) is implemented, consisting of multiple upsampling steps to decode the hidden features and output the final segmentation mask. After reshaping the sequence of hidden features $\bm{\mathrm{z}}_{L}\in\mathbb{R}^{\frac{HW}{P^{2}}\times D}$ to the shape $\frac{H}{P}\times\frac{W}{P}\times D$, CUP is instantiated by cascading multiple upsampling blocks to reach the full resolution from $\frac{H}{P}\times\frac{W}{P}$ to $H\times W$. Each block consists of a 2×2$\times$2 × upsampling operator, a $3\times 3$ convolution layer, and a ReLU layer successively. CUP, together with the hybrid encoder, forms a U-shaped architecture that enables feature aggregation at different resolution levels via skip connections. The upsampling process is also illustrated in Figure 2.
3.2 Robust Cell Segmentation via Virtual Outlier Sampling
To enhance the robustness of our model and accurately simulate the different stages of pluripotent carcinoma cells, we propose the sampling of virtual outliers during the training process. Traditional deep learning models for image segmentation typically utilize CNN frameworks, which compress the latent space into lower dimensions through feature extraction. In contrast, our model is built on a Vision Transformer architecture, which poses challenges in synthesizing virtual outliers directly in the high-dimensional pixel space due to its complexity and the optimization difficulties involved. Despite the CNN-Transformer hybrid structure as its encoder in the form of TransUNet, the latent features of high-dimensional images remain in a large matrix. To address this, we opt to sample directly from the output logits of the segmentation map, denoted by $s(\mathbf{x}_{i})$, instead of sampling from the latent space.
Given the generic nature of our method, we establish general notations applicable to similar tasks in image classification or segmentation. We model the features of all target classes as conditional multivariate Gaussian distributions:
$p_{\theta}(s(\mathbf{x})|y=k)=\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}),$ (6)
where $\boldsymbol{\mu}_{k}$ represents the mean for class $k\in\{1,2,\dots,K\}$, and $\boldsymbol{\Sigma}$ is the covariance matrix shared across classes. The function $s(\mathbf{x})\in\mathbb{R}^{m}$ represents the segmentation map of a cell instance $\mathbf{x}$, where $m$ is significantly smaller than the input dimension $d$. In our specific application to the pluripotent carcinoma cell dataset, we focus on two classes: the cells and the background, implying $k=2$.
To estimate the parameters of the Gaussian distribution, we calculate the empirical class means and covariance matrix from the training samples $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$:
$\widehat{\boldsymbol{\mu}}_{k}$ $=\frac{1}{N_{k}}\sum_{i:y_{i}=k}s(\mathbf{x}_{i}),$ (7)
$\widehat{\boldsymbol{\Sigma}}$ $=\frac{1}{N}\sum_{k=1}^{K}\sum_{i:y_{i}=k}\left(s(\mathbf{x}_{i})-\widehat{\boldsymbol{\mu}}_{k}\right)\left(s(\mathbf{x}_{i})-\widehat{\boldsymbol{\mu}}_{k}\right)^{\top}.$ (8)
where $N_{k}$ is the number of objects in class $k$, and $N$ is the total number of objects. We use online estimation for efficient training, where we maintain a class-conditional queue with $|Q_{k}|$ object instances from each class. In each iteration, we enqueue the embeddings of objects to their corresponding class-conditional queues and dequeue the same number of object embeddings.
Sampling from the Feature Representation Space We propose sampling virtual outliers from the feature representation space of cell images, utilizing the multivariate distributions estimated from these features. These virtual outliers are designed to refine the decision boundary between different stages of cell development and other outlier data, which may not be adequately represented in the training set.
To accomplish this, we generate virtual outliers $\mathcal{V}_{k}$ from the $\epsilon$-likelihood region of the estimated class-conditional distribution for each cell stage:
$\mathcal{V}_{k}=\left\{v_{k}\mid\frac{1}{(2\pi)^{m/2}|\widehat{\Sigma}|^{1/2}}\exp\left(-\frac{1}{2}(v_{k}-\widehat{\mu}_{k})^{\top}\widehat{\Sigma}^{-1}(v_{k}-\widehat{\mu}_{k})\right)<\epsilon\right\},$ (9)
where $v_{k}\sim\mathcal{N}(\widehat{\mu}_{k},\widehat{\Sigma}_{k})$ represents the sampled virtual outliers for class $k$. These samples are chosen from a sublevel set defined by a small $\epsilon$, ensuring that the outliers are near the decision boundary of the class. We hypothesize that most of the cell pixels are near the boundary of existing cells as shown in Figure 3. Compared to the ground truth masks shown in the second subfigure in Figure 3, the synthesized outliers contain more information about the cells, which are difficult for humans to annotate due to their small size. This strategy aims to enhance the model’s sensitivity to subtle yet critical variations in cell morphology that might indicate different developmental stages or pathological conditions.
(a) Original cell images from the training set
(b) Corresponding ground truth masks
(c) Visualization of the proposed segmentation map with synthesis outliers
Figure 3: Examples showcasing the training with visual outliers. Patches are taken from the training set, presented in sizes of $224\times 224$ and $448\times 448$. Compared to the ground truth masks, the synthetic outlier samples simulate small cell dots at the pixel level, which are challenging for human annotation. Additionally, small white dots are presented as outliers, which may be difficult to discern without magnification.
3.3 Unknown-aware Training Objective
The objective is to develop an uncertainty-aware deep learning model capable of predicting a segmentation map $\hat{S}$ for input cell images. To this end, we introduce a new training objective for unknown-aware learning, which utilizes virtual outliers as discussed in Section 3.2. We start by defining the standard Dice loss between the predicted pixel scores $\hat{y}$ and the ground truth targets $y$:
$\mathcal{L}_{\text{Dice}}(f(\mathbf{x};\theta),y)=1-\frac{2\sum_{i}p_{i}g_{i}+\epsilon}{\sum_{i}p_{i}^{2}+\sum_{i}g_{i}^{2}+\epsilon},$ (10)
where $p_{i}$ represents the predicted probabilities for the actual class labels $g_{i}$, and $\epsilon$ is a small constant introduced to prevent division by zero.
In the case of medical segmentation [51], Dice loss is frequently combined with Cross Entropy loss to form the overall training objective. The Cross Entropy loss is defined as follows:
$\mathcal{L}_{\text{CE}}(f(\mathbf{x};\theta),y)$ $=-\sum_{i}y_{i}\log(f(\mathbf{x}_{i};\theta)),$ (11)
where $f(\mathbf{x}_{i};\theta)$ is the model’s output for the input $\mathbf{x}_{i}$, and $y_{i}$ is the corresponding true label.
The overall training objective for medical segmentation, as used in vanilla TransUNet, combines both Cross Entropy and Dice losses [52]. This combined loss function is employed as the primary training loss, defined as:
$\mathcal{L}_{\text{seg}}=\lambda_{1}\mathcal{L}_{\text{CE}}(\hat{S}_{l},D(S))+\lambda_{2}\mathcal{L}_{\text{DICE}}(\hat{S}_{l},D(S)).$ (12)
The weights $\lambda_{1}$ and $\lambda_{2}$ are used to balance these two terms, typically set to be equally distributed [1].
With the virtual outliers $v_{k}$ sampled from the predicted segmentation map $s(\mathbf{x}_{i})$, we effectively capture the inherent uncertainty of the segmentation process. To make our model uncertainty-aware, we adapt both the Dice loss and Cross Entropy loss for training as follows:
$\mathcal{L}_{\text{Dice\_out}}(f_{\text{syn}}(\mathbf{x};\theta),y)$ $=1-\frac{2\sum_{i}p_{\text{syn},i}g_{i}+\epsilon}{\sum_{i}p_{\text{syn},i}^{2}+\sum_{i}g_{i}^{2}+\epsilon},$ (13)
$\mathcal{L}_{\text{CE\_out}}(f_{\text{out}}(\mathbf{x};\theta),y)$ $=-\sum_{i}y_{i}\log(f_{\text{out}}(\mathbf{x}_{i};\theta)),$ (14)
where $p_{\text{syn},i}$ represents the predicted probabilities for the actual class labels $g_{i}$ from the synthetic outputs, and $f_{\text{out}}(\mathbf{x}_{i};\theta)$ denotes the model’s output for synthetic data designed to simulate under-represented features.
Furthermore, we model the uncertainty captured by the virtual outliers through a specific loss function:
$\mathcal{L}_{\text{uncertainty}}=\beta_{1}\mathcal{L}_{\text{CE\_out}}(\hat{S}_{l},D(S))+\beta_{2}\mathcal{L}_{\text{DICE\_out}}(\hat{S}_{l},D(S)),$ (15)
where $\beta_{1}$ and $\beta_{2}$ are coefficients that weigh the contributions of the cross-entropy and Dice losses, respectively, in quantifying the uncertainty associated with the segmentation of cells.
Overall Training Objective. Although the proposed loss function is initially designed for binary segmentation, it is adaptable for multi-class segmentation scenarios. Our overall training objective integrates the segmentation loss, which includes both Dice and Cross Entropy losses, and a regularization component that addresses model uncertainty:
$\min_{\theta}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\lambda\mathcal{L}_{\text{seg}}(\mathbf{x},y;\theta)+\beta\mathcal{L}_{\text{uncertainty}}(\mathbf{x},y;\theta)\right],$ (16)
where $\lambda$ and $\beta$ is the weighting coefficient for the uncertainty regularization. The segmentation loss, $\mathcal{L}_{\text{seg}}$, combines the Dice loss, $\mathcal{L}_{\text{DICE}}$, and the cross-entropy loss, $\mathcal{L}_{\text{CE}}$, formulated to effectively handle the segmentation of cells across different stages and morphologies. The uncertainty component, $\mathcal{L}_{\text{uncertainty}}$, leverages synthetic data to enhance the model’s ability to predict underrepresented features and quantify predictive confidence. Ablation studies, detailed in Section 5.4, demonstrate the efficacy of this loss function in improving segmentation accuracy and model robustness.
4 Dataset Analysis
4.1 Data Description
[…]
The NT2 dataset consists of 105 high-resolution raw images of NT2 cells undergoing differentiation, of which 36 images are fully annotated. These images have resolutions primarily around $4000\times 3000$ pixels, enabling detailed morphological analysis and accurate segmentation.
4.2 Data Preprocessing
Data Augmentation. To efficiently train our TransUNet model, we have devised a method to augment the dataset by transforming the high-resolution images into smaller, manageable patches. This is achieved by dividing the images into patches of sizes $224\times 224$, $448\times 448$, $1000\times 1000$, $1500\times 1500$, and $2000\times 2000$ pixels. We employ an overlapping strategy with an overlap ratio of 0.35 to preserve the spatial context across adjacent patches, which is crucial for capturing the comprehensive spatial information of the cell structures. Combined with the original undivided images and masks, this makes the full dataset for training and testing.
Using this technique, we generate a total of 15,710 patches. These patches are then randomly split into a training set and a testing set, maintaining an 80% to 20% ratio, respectively. This distribution results in 12,568 patches for training and 3,142 patches for testing. To optimize the model’s performance, we implement a data pruning strategy during both the training and testing stages, excluding all patches that do not contain any part of the annotated masks, which will reduce the training set to 10,016 samples for testing.
5 Experiments
5.1 Implementation details
We evaluate our model against established benchmarks including Cellpose and TransUNet, utilizing the AdamW optimizer with exponential learning rate decay on the R50-ViT-B architecture. Specifically for Cellpose, we utilize two of the highest performing pre-trained configurations, cyto3 and cyto2, both set with a cell diameter of $30$. These models are trained with a learning rate of $0.01$, a weight decay of $5\times 10^{-5}$, and over $200$ epochs. The batch size for training is set at $8$, with verbose output enabled for detailed progress tracking. During testing, we adapted the Cellpose model to handle larger cells by setting the diameter parameter to $100$, with no flow threshold and a cell probability threshold of $0.2$, adjusting the channels accordingly to match specific imaging conditions. For the TransUNet, we adhere which includes two types of encoders: a pure Transformer-based encoder, denoted as “ViT-B16,” and a hybrid encoder that combines ResNet-50 with ViT, denoted as “R50-B16”. The input resolution and patch size $P$ are set to 224$\times$224 and 16, respectively. We set the learning rate to 0.01, momentum to 0.9, and weight decay to $1\times 10^{-4}$. The models are trained over 200 epochs, with sampling commencing at the 150th epoch. Both models are trained on two NVIDIA A100 GPUs with a batch size of 256, focusing on 10,000 critical samples by selectively sampling 1,000 pixels per image to enhance training efficiency and effectiveness.
5.2 Evaluation Metrics
To evaluate our method, we considered four key metrics: 1) the average Dice Similarity Coefficient (DSC) [53], 2) the average Hausdorff Distance (HD) [54], 3) the Intersection over Union (IoU) [55], and 4) the mean Average Precision (mAP) [56]. These metrics provide a comprehensive assessment of segmentation accuracy and model performance.
For each class, the unique matching of $(y,\hat{y})$ splits the predicted and the GT segments into three sets:
  • True Positives (TP): the number of correctly predicted pixels that match the ground truth.
  • False Positives (FP): the number of pixels incorrectly predicted as part of the object when they are not in the ground truth.
  • False negatives (FN): the number of pixels that are missed by the prediction but are present in the ground truth.
The Dice Similarity Coefficient (DSC) is a standard metric for assessing the overlap between the predicted segmentation mask and the ground truth mask. It is defined as:
DSC $=\frac{2\times\text{TP}}{2\times\text{TP}+\text{FP}+\text{FN}},$ (17)
where TP, FP, and FN denote the number of true positive, false positive, and false negative pixels, respectively. The Mean DSC is computed by averaging the DSC values across all segmentation instances or classes in the dataset. The Mean DSC provides a robust measure of overall segmentation accuracy, facilitating a comprehensive evaluation of model performance across all classes or instances. Since we only have $1$ target class which is the cell, we only use the DSC as the evaluation metric.
To further assess the boundary accuracy, we employ the 95th Percentile Hausdorff Distance (HD95), which offers a more robust measure of the spatial discrepancy between the predicted and ground truth segmentation boundaries. The Hausdorff Distance (HD) between two sets, $X$ and $Y$, is given by:
$\text{HD}(X,Y)$ $=\max(h(X,Y),h(Y,X)),$ (18)
where
$h(X,Y)=\max_{x\in X}\min_{y\in Y}d(x,y),$ (19)
and $d(x,y)$ denotes the Euclidean distance between points $x\in X$ and $y\in Y$.
The HD95 is defined as the 95th percentile of the sorted set of all distances between the predicted and ground truth boundaries:
[…]
where $d(x,Y)=\min_{y\in Y}d(x,y)$ and $d(y,X)=\min_{x\in X}d(y,x)$. The HD95 metric reduces the influence of outliers by focusing on the 95th percentile distance, providing a robust measure of boundary alignment between the predicted and ground truth segmentations.
The model evaluation should not only use the statistic metrics but also need a qualitative evaluation from the actual segmentation view. The intersection over union (IoU) [57] as an important evaluation index in semantic segmentation measures the overlap of the ground truth and prediction region. It is also calculated:
IoU=Ground TruthPredictionGround $\text{IoU}=\frac{\text{Ground Truth}\bigcap\text{Prediction}}{\text{Ground Truth}\bigcup\text{Prediction}},$ (21)
where $\bigcap$ denotes the intersection operator and $\bigcup$ represents the union operator. The IoU is generally calculated based on categories, which is to accumulate the IoU value of each category. The IoU value is to average the sum IoU results of each category to obtain a global evaluation. Therefore, the IoU is actually the mean value, that is, the average crossover ratio (mean IoU).
Finally, the mean Average Precision (mAP) is defined as:
mAP $=\frac{1}{N}\sum_{i=1}^{N}\text{AP}_{i},$ (22)
where $\text{AP}_{i}$ represents the average precision for the $i$-th class, and $N$ is the total number of classes.
These metrics collectively enable a robust evaluation of our model’s performance in terms of segmentation accuracy and overall effectiveness.
5.3 Comparison with Baselines
Table. 1 presents the main results of the proposed approach, which includes training with outliers compared to the Cellpose pretrained on cyto3 and cyto2, and TransUNet of ViT-B16 and R50-B16 variants. Despite the advantages of the Cellpose model, such as requiring a smaller number of annotations, it does not offer superior performance due to its reliance on the flow and diameter of cells. In our study, these characteristic properties vary across different stages, and the irregular shapes of cells do not provide consistent diameters for effective segmentation. In comparison, our training approach with Outlier Synthesis enhances the performance significantly over the state-of-the-art TransUNet across all types of loss combinations.
Method DSC HD95 IoU-0.5 IoU-0.75 IoU-0.9 mAP
Framework Encoder Loss $\uparrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
Cellpose [6] cyto3 63.66 137.9 25.31 12.35 3.641 48.51
cyto2 48.15 181.5 19.02 8.220 2.569 36.78
TransUnet [1] ViT-B16 72.47 38.21 76.42 30.19 3.019 36.54
R50-B16 80.35 21.98 88.00 58.67 9.220 51.97
MorphoSeg ViT-B16 Balance 72.78 38.69 77.35 30.19 3.787 37.46
Norm 74.24 36.82 81.54 43.82 4.765 38.96
Pareto 73.67 37.10 80.43 42.66 3.433 37.58
R50-B16 Balance 86.33 16.33 93.02 79.68 20.11 64.27
Norm 84.40 26.37 91.64 72.50 11.30 58.48
Pareto 86.57 15.75 93.02 80.01 21.87 64.96
Table 1: Comparison on our proposed cell dataset. All methods incorporating Outlier Synthesis are trained with a sample size of $100,000$, and the selection count is set to $10,000$. $\uparrow$ indicates that larger values are preferable, while $\downarrow$ indicates that smaller values are better. All numerical values are presented as percentages. Bold numbers indicate superior results.
5.4 Ablation Studies
Incorporating Uncertainty into Loss Function Design. To determine the optimal values for the weight parameters $\lambda$ and $\beta$ in Equation.(16), we evaluate three sophisticated approaches to weight the loss components, tailored for enhanced model training and testing. The Original Balancing method, consistent with TranUNet, evenly distributes weights across all loss components, simplifying optimization but potentially overlooking specific task requirements within the model. Alternatively, the Normalized Losses method dynamically scales each loss component by its magnitude:
$\mathcal{L}_{\text{norm}}=\sum_{i}\frac{\mathcal{L}_{i}}{\|\mathcal{L}_{i}\|}+\epsilon,$ (23)
[…]
$\mathcal{L}_{\text{pareto}}=\mathcal{L}_{\text{loss}}+\sum_{i\neq\text{loss}}\frac{\mathcal{L}_{i}}{\|\mathcal{L}_{i}/\mathcal{L}_{\text{loss}}\|},$ (24)
which potentially provides a more rational scaling by relating all losses back to the primary loss component. Lastly, the Pareto Optimization strategy [58], aims to achieve a Pareto efficient solution by treating each loss as an objective in a multi-objective optimization framework, wherein improvements in one objective are sought without significant compromises in others, facilitating an equilibrium that optimizes overall model performance.
Figure 4: Visualization of Ntera-2 Cell Segmentation in the Testing Dataset: Original Images, Segmentation Masks, and Overlays. Row A displays the original images, Row B displays the segmentation mask and Row C displays the mask overlay on the original image. Image set 1 highlights the ability of the network to segment cells in an unfocused image. Image set 2 highlights the ability of the network to segment cells despite low contrast and unclear cell borders. Images 3 and 4 display ability of the network to segment high-quality images at different magnifications, x10 and x25, respectively.
5.5 Patch-Based Approach for High-Resolution Inference and Visualization
Given the inherent limitations of Vision Transformers, where token size constraints can result in vague and imprecise inference on large images, we propose an advanced patch-based inference technique to address the sensitivity to pixel-level cellular details. This approach enhances segmentation accuracy for high-resolution biological cell images while making the framework more practical and accessible for researchers. Example visualizations of the results are presented in Figure 4.
During a single image inference stage, the image is initially subdivided into patches of size $224\times 224$, similar to the data augmentation strategy presented in Section 4.2. This subdivision allows the model to focus on and accurately detect fine-grained variations at the boundaries of the Ntera-2 cells and their surrounding environment. Each patch undergoes a separate inference process. To preserve the global spatial context, we incorporate an overlapping strategy with a $56$-pixel margin. This overlap ensures continuity and minimizes edge artifacts. The inference results from the patches are then combined using an average pooling strategy and visualized as segmentation masks, shown in red in Figure 4. This end-to-end training and visualization strategy significantly facilitates the application of our framework, enabling researchers to use it without requiring further modifications.
6 Conclusion and Future Work
In this paper, we introduced a novel data repository containing differentiating NT2 cells, designed to facilitate the segmentation of complex cell morphologies. This dataset requires pixel-level segmentation, which presents new challenges for accurately segmenting cells with intricate shapes. To address these challenges, we proposed an uncertainty-aware segmentation method MorphoSeg enhancing the base TranUNet by training on virtual outliers sampled near the feature representation space, thereby increasing the model’s sensitivity to small pixel variations and morphological changes across different cell stages. This enhancement makes our MorphoSeg more robust and effective for segmenting complex and irregularly shaped cells.
The experimental results demonstrate that the proposed MorphoSeg outperforms existing benchmarks, including Cellpose and TransUNet. Our best-performing model, MorphoSeg with a hybrid encoder, showed a notable improvement in performance metrics, achieving a 6.23% increase in the Dice Similarity Coefficient (DSC). This statistic measures the similarity between the segmented cells and the ground truth. A reduction in HD95 (Hausdorff Distance at the 95th percentile) was also measured, indicating a decreased spatial discrepancy between the predicted and actual segmentation boundaries compared to the baseline configuration.
By providing comprehensive visualizations and complete code, we aim to contribute to not only advance further research in this emerging domain but also to serve as a practical solution for cell segmentation and visualization in diverse medical AI applications.
Future Work: Building on the proposed NT2 cells dataset, future research will address the challenge of segmenting complex cell morphologies, particularly those with substantial overlapping regions, which remain a significant obstacle to achieving accurate segmentation. Further efforts will focus on developing advanced deep-learning methods for biological AI to overcome these challenges.
Acknowledgments
We express our gratitude to the UK Engineering and Physical Sciences Research Council (EPSRC) for their support through the following projects: EP/V026747/1 (Trustworthy Autonomous Systems Node in Resilience) and EP/T013265/1 (NSF-EPSRC: ShiRAS. Towards Safe and Reliable Autonomy in Sensor Driven Systems). We also acknowledge the support of the National Science Foundation (NSF) under Grant No. ECCS 1903466 for the ShiRAS project. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript (AAM) version arising from this work.